10.Use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Just determine the number and types of solutions.



2x2 - 6x + 5 = 0

Please help. Thank you.

Answer by uday1435(57)

2x2 - 6x + 5 = 0
Comparing with the general equation ax^2 + bx + c =0
a = 2 , b = -6 and c = 5
The discriminant b^2 – 4ac = 36 – 4(2)(5) = 36 - 40 = - 4
Ans: Since the discriminant is negative there are two complex roots




If you still require further clarifications please write to udayakumar.t.r@gmail.com